OpenStudy (anonymous):

What is the difference in the y-values between(1,7) and (3,11? what is half of that difference?how can you use that different to find the y-value of the midpoint?

OpenStudy (anonymous):

difference is (11-7) = 4 add half the difference to the lower value to get mid-point y y = 7 + 4/2 = 9